40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  SUS Lab eLog  Not logged in ELOG logo
Entry  Wed Jul 7 13:05:07 2021, Paco, DailyProgress, 1418 nm AUX ECDL, New aspheric flexures 
    Reply  Fri Jul 9 11:40:42 2021, Paco, DailyProgress, 1418 nm AUX ECDL, New aspheric flexures IMG_1773.jpeg
       Reply  Wed Jul 14 12:21:08 2021, Paco, DailyProgress, 1418 nm AUX ECDL, ECDL lases... and MZ locked mid-fringe selfhomodyne_1418.jpg
          Reply  Wed Aug 4 11:36:30 2021, Radhika, DailyProgress, 1418 nm AUX ECDL, 1419 nm ECDL with 2um AOM tests rf_setup.jpgdiffraction_levels.png
             Reply  Tue Aug 10 09:51:44 2021, Paco, DailyProgress, 1418 nm AUX ECDL, 1419 nm ECDL AOM diffraction at 95 MHz rf_setup.jpg
                Reply  Thu Aug 12 11:49:59 2021, Radhika, DailyProgress, 1418 nm AUX ECDL, 1419 nm ECDL AOM diffraction at 95 MHz zeroth_order_contour.pdffirst_order_contour.pdf
                   Reply  Tue Aug 17 11:09:29 2021, rana, DailyProgress, 1418 nm AUX ECDL, 1419 nm ECDL AOM diffraction at 95 MHz 
                   Reply  Wed Sep 1 13:12:02 2021, Radhika, DailyProgress, 1418 nm AUX ECDL, 1419 nm ECDL AOM diffraction at 95 MHz 
Message ID: 1922     Entry time: Wed Sep 1 13:12:02 2021     In reply to: 1920
Author: Radhika 
Type: DailyProgress 
Category: 1418 nm AUX ECDL 
Subject: 1419 nm ECDL AOM diffraction at 95 MHz 

[Paco, Radhika]

Today we tried to pick up from [1920] by repeating the sweep measurements across RF frequency, at 3 dBm (max power). We noticed that the 0th order signal would dip around the expected value, consistent with the plot in [1920]. However, there was no signal from the 1st order. Clearly diffraction was occurring as seen by the dip in 0th order, but nothing was coming out of the 1st order port. We spent some time debugging by swapping the photodetector inputs / playing with the PD gains / performing power cycles, but got no insight into the issue. 

We suspected the 1st order fiber coming out of the AOM might be damaged, since it loops around fairly tightly. After giving it more slack, we still saw no signal. We wanted to test the fiber, so we took an unused output of the 50-50 beamsplitter and fed it into the 1st order port, effectively running the AOM in reverse. We hooked up the input and 0th order ports to the photodiodes and did not observe any signal. From here we were more convinced that the 1st order fiber may have seen some damage. 

For next steps, we can still use the existing fiber setup to take measurements of relative intensity noise (RIN), using the 0th order output of the AOM. I plan to do this in the next few days. Meanwhile, Paco is looking into ordering parts for a free space setup. We found a free-space AOM at 1064nm that seems promising, and we will work to transition the setup accordingly. 

ELOG V3.1.3-